# Convergence and long-time behavior of finite volumes for a generalized Poisson–Nernst–Planck system with cross-diffusion and size exclusion

### v2

@@ 1.1. The continuous generalized Poisson–Nernst–Planck model
$\mathcal{A}=\left\{U=\left(u_{i}\right)_{1\leq i\leq I}\in({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}})^{I}\;\middle|\;\sum_{i=1}^{I}u_{i}\leq 1\right\}$
@@ 1.3. Weak solution
(12) $0\leq u_{i}\leq 1$ for $i=0,\dots,I$ and a.e. $(t,x)\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega$.
@@
(13) $\|\phi\|_{L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)}+\|\phi-\phi^{D}\|_{L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};V)}\leq C(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(z_{i})_{i}},\Omega,f),$
@@
[…]
(14) $\iint_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega}\sum_{i=1}^{I}D_{i}u_{0}^{2}e^{-z_{i}\phi}|\nabla\sqrt{w_{i}}|^{2}\leq C.$
As already noticed in [31], the above inequality yields a control in $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};H^{1}(\Omega))$ on $\sqrt{u_{0}}$ and on $u_{0}$ itself, as well as control on some product terms. More precisely, in the numerical analysis exposed hereafter, we derive a $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};H^{1}(\Omega))$ estimate on $u_{i}\sqrt{u_{0}}$, which, together with (12) provides a $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};H^{1}(\Omega))$ control on $u_{i}u_{0}$ too. Hence, all the terms in the following expression of the fluxes
@@
[…]
  • $\phi\in L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$ with $\phi-\phi^{D}\in L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};V)$ satisfies
@@
  • for all $\varphi\in C^{1}_{c}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})$ and all $i=1,\dots,I$, there holds
    (17) $\iint_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega}u_{i}\partial_{t}\varphi+\int_{\Omega}u_{i}^{0}\varphi(0,\cdot)-\iint_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega}D_{i}\big{(}\nabla(u_{0}u_{i})-4u_{i}\sqrt{u_{0}}\nabla\sqrt{u_{0}}+u_{i}u_{0}z_{i}\nabla\phi\big{)}\cdot\nabla\varphi=0.$
@@ 1.4. Equilibrium and modified Poisson–Boltzmann equation
$\partial_{y}r(y,\boldsymbol{\xi})=\sum_{i=0}^{I}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z_{i}^{2}}v_{i}(y,\boldsymbol{\xi})-\left(\sum_{i=0}^{I}z_{i}v_{i}(y,\boldsymbol{\xi})\right)^{2}\geq 0,$
@@
One readily checks that $\Psi$ is strictly convex, and thus the minimiser $(\phi^{\infty},\boldsymbol{\mu}^{\infty})$ is unique. Moreover, one easily verifies that minimizers for $\Psi$ are steady states of the system (1)–(7). Despite hints in these directions, the uniqueness of the steady state is neither established in [6], nor, up to our knowledge, in more recent contributions. Furthermore, no quantitative convergence result has been established so far. Non-quantitative results were however recently presented in [19] in the absence of electrical interactions, i.e. $z_{i}=0$ for all $i$. This gap is partially filled in the discrete case. We rigorously establish in Section 5 the uniqueness of the steady states, which are minimisers of some discrete counterpart of the functional $\Psi$, cf (35), and the (non-quantitative) convergence towards such a steady profile as time goes to $\infty$.
@@ 1.5. Goal and positioning of the paper
We also investigate the long-time behaviour of the numerical scheme. We show in particular that the discrete solution to our scheme for the evolution problem converges as time goes to infinity towards the unique solution to the discrete steady problem. As the proof relies on LaSalle’s invariance principle, no quantitative convergence rate is derived, similarly to what is proposed in [19] in the continuous setting but in absence of electrical interactions. By establishing this result, we rigorously establish in the discrete setting the long-time behavior conjectured in [6] and recalled in Section 1.4. Numerical simulations suggest that the convergence rate strongly depends on the Debye length, but also on the initial profile $U^{0}$ (in opposition to the classical Poisson–Nernst–Planck problem without volume filling [3]).
@@ 2.1. Space and time discretizations
For any $K\in\mathcal{T}$, we assume that there exists a subset $\mathcal{E}_{K}$ of distinct elements of $\mathcal{E}$ such that the boundary of a control volume can be described by $\partial K=\bigcup_{\sigma\in\mathcal{E}_{K}}\sigma$ and, consequently, $\mathcal{E}=\bigcup_{K\in\mathcal{T}}\mathcal{E}_{K}$. Additionally, we assume that boundary edges $\mathcal{E}_{\text{ext}}=\mathcal{E}\setminus\mathcal{E}_{\rm int}$ are either subsets of $\Gamma^{D}$ or $\Gamma^{N}$. To each control volume $K\in\mathcal{T}$ we assign a cell center ${x}_{K}\in K$ which satisfies the orthogonality condition: If $K,L$ share a face $\sigma=K|L$, then the vector $x_{L}-x_{K}$ is orthogonal to $\sigma=K|L$.The triplet $\left(\mathcal{T},\mathcal{E},\{{x}_{K}\}_{K\in\mathcal{T}}\right)$ is called an admissible mesh.
@@
[…]
@@ 2.3. Main results and organization of the paper
$\phi_{\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}\phi\quad\text{and}\quad u_{0,\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}u_{0}\quad\text{in}\;L^{p}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};L^{p}(\Omega))\quad\forall p\in[1,+\infty),$
@@
$U_{\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}U\quad\text{in the }L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)^{I}\text{ weak-$\star$ sense}.$
@@ 3.1. Uniform a priori bounds and existence of a discrete solution
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}}\;:\;\begin{cases}[0,\tau^{n}]\times[-2,2]^{I\times\mathcal{T}}\times[-C-1,C+1]^{\mathcal{T}}\to\mathbb{R}^{\mathcal{T}}\times\mathbb{R}^{\mathcal{T}}\\ (s,U,\phi)\mapsto{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}}(s,U,\phi)\end{cases}$
@@
Thus, the topological degree corresponding to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}}(s,U,\phi)=0$ is equal to one, all along the homotopy, and hence in particular for $s=\tau^{n}$. That implies the existence of a solution to the scheme (but it does not guarantee uniqueness). ∎
@@ 3.2. Energy dissipation at the discrete level
(41) $\mathcal{H}_{\mathcal{T}}^{n}=\sum_{K\in\mathcal{T}}m_{K}H(U_{K}^{n})+\frac{\lambda^{2}}{2}\sum_{\sigma\in\mathcal{E}}a_{\sigma}(\phi_{K}^{n}-\phi_{K\sigma}^{n})^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\lambda^{2}\sum_{\sigma\in\mathcal{E}^{D}}a_{\sigma}\phi_{\sigma}^{D}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\phi_{\sigma}^{D}-\phi_{K}^{n})},$
@@
where $\widetilde{H}:\left\{(u_{i})_{0\leq i\leq I}\;\middle|\;\sum_{i=0}^{I}u_{i}=1\right\}\to[-\log(I+1),+\infty)$, defined by $\widetilde{H}(u_{0},\dots,u_{I}):=\sum_{i=0}^{I}\log(u_{i})$, and
@@ 3.3. Further uniform estimates on the discrete solution
$\mathcal{D}_{i,\sigma}^{n}\geq\kappa a_{\sigma}\cosh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right)\left(\sqrt{u_{i,K}^{n}u_{0,L}^{n}}-\sqrt{u_{i,L}^{n}u_{0,K}^{n}}\right)^{2}\\ -{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\kappa a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,L}^{n}}+\sqrt{u_{i,L}^{n}u_{0,K}^{n}}\right)^{2}\sinh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right).$
@@
(55) $\sum_{n=1}^{N}\sum_{K\in\mathcal{T}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m_{K}}(u_{i,K}^{n}-u_{i,K}^{n-1})\varphi_{K}^{n}\leq C\left(1+t^{N}\right)^{1/2}\left(\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\varphi_{K}^{n}-\varphi_{K\sigma}^{n}\right)^{2}\right)^{1/2}.$

### v3

# arXiv:2411.11583v3[math.NA] 03 Mar 2026

@@ 1.1. The continuous generalized Poisson–Nernst–Planck model
$\mathcal{A}=\left\{U=\left(u_{i}\right)_{1\leq i\leq I}\in(\mathbb{R}_{\geq 0})^{I}\;\middle|\;\sum_{i=1}^{I}u_{i}\leq 1\right\}$
@@ 1.3. Weak solution
(12) $0\leq u_{i}\leq 1$ for $i=0,\dots,I$ and a.e. $(t,x)\in\mathbb{R}_{\geq 0}\times\Omega$.
@@
(13) $\|\phi\|_{L^{\infty}(\mathbb{R}_{\geq 0}\times\Omega)}+\|\phi-\phi^{D}\|_{L^{\infty}(\mathbb{R}_{\geq 0};V)}\leq C(\lambda,(z_{i})_{i},\Omega,f),$
@@
is equipped with the $H^{1}(\Omega)$ semi-norm $\|\nabla v\|_{L^{2}}$, which is a norm thanks to the Poincaré inequality. We have emphasized in the right-hand side of (13) the dependence of the bound on the data, especially on the Debye length $\lambda$. The estimate (9) is the other main estimate on which our study will rely. It provides enough control to define a proper notion of weak solutions, but at the price of a suitable reformulation of the fluxes (2). Let us first remark that, since the free energy $\mathcal{H}$ is bounded (see Lemma 3.7), integrating (9) over $t\in\mathbb{R}_{\geq 0}$ provides
(14) $\iint_{\mathbb{R}_{\geq 0}\times\Omega}\sum_{i=1}^{I}D_{i}u_{0}^{2}e^{-z_{i}\phi}|\nabla\sqrt{w_{i}}|^{2}\leq C.$
As already noticed in [31], the above inequality yields a control in $L^{2}_{\rm loc}(\mathbb{R}_{\geq 0};H^{1}(\Omega))$ on $\sqrt{u_{0}}$ and on $u_{0}$ itself, as well as control on some product terms. More precisely, in the numerical analysis exposed hereafter, we derive a $L^{2}_{\rm loc}(\mathbb{R}_{\geq 0};H^{1}(\Omega))$ estimate on $u_{i}\sqrt{u_{0}}$, which, together with (12) provides a $L^{2}_{\rm loc}(\mathbb{R}_{\geq 0};H^{1}(\Omega))$ control on $u_{i}u_{0}$ too. Hence, all the terms in the following expression of the fluxes
@@
  • $u_{i}\in L^{\infty}(\mathbb{R}_{\geq 0}\times\Omega)$ with $U\in\mathcal{A}$ a.e. in $\mathbb{R}_{\geq 0}\times\Omega$, with moreover $\sqrt{u_{0}}$ and $u_{i}\sqrt{u_{0}}$ belonging to $L^{2}_{\rm loc}(\mathbb{R}_{\geq 0};H^{1}(\Omega))$ for $u_{0}$ defined as in (3);
  • $\phi\in L^{\infty}(\mathbb{R}_{\geq 0}\times\Omega)$ with $\phi-\phi^{D}\in L^{\infty}(\mathbb{R}_{\geq 0};V)$ satisfies
@@
  • for all $\varphi\in C^{1}_{c}(\mathbb{R}_{\geq 0}\times\overline{\Omega})$ and all $i=1,\dots,I$, there holds
    (17) $\iint_{\mathbb{R}_{\geq 0}\times\Omega}u_{i}\partial_{t}\varphi+\int_{\Omega}u_{i}^{0}\varphi(0,\cdot)-\iint_{\mathbb{R}_{\geq 0}\times\Omega}D_{i}\big(\nabla(u_{0}u_{i})-4u_{i}\sqrt{u_{0}}\nabla\sqrt{u_{0}}+u_{i}u_{0}z_{i}\nabla\phi\big)\cdot\nabla\varphi=0.$
@@ 1.4. Equilibrium and modified Poisson–Boltzmann equation
$\partial_{y}r(y,\boldsymbol{\xi})=\sum_{i=0}^{I}z_{i}^{2}v_{i}(y,\boldsymbol{\xi})-\left(\sum_{i=0}^{I}z_{i}v_{i}(y,\boldsymbol{\xi})\right)^{2}\geq 0,$
@@
One readily checks that $\Psi$ is strictly convex, and thus the minimiser $(\phi^{\infty},\boldsymbol{\mu}^{\infty})$ is unique. Moreover, one easily verifies that minimizers for $\Psi$ are steady states of the system (1)–(7). Despite hints in these directions, the uniqueness of the steady state is neither established in [6], nor, to our knowledge, in more recent contributions. Furthermore, no quantitative convergence result has been established so far. Non-quantitative results were however recently presented in [19] in the absence of electrical interactions, i.e. $z_{i}=0$ for all $i$. We partially fill this gap in the discrete case. We rigorously establish in Section 5 the uniqueness of the steady states, which are minimisers of some discrete counterpart of the functional $\Psi$, cf. (35), and the (non-quantitative) convergence towards such a steady profile as time goes to $\infty$.
@@ 1.5. Goal and positioning of the paper
We also investigate the long-time behaviour of the numerical scheme. We show in particular that the discrete solution to our scheme for the evolution problem converges as time goes to infinity towards the unique solution to the discrete steady problem. As the proof relies on LaSalle’s invariance principle, no quantitative convergence rate is derived, similarly to what is proposed in [19] in the continuous setting but in the absence of electrical interactions. By establishing this result, we rigorously establish in the discrete setting the long-time behavior conjectured in [6] and recalled in Section 1.4. Numerical simulations suggest that the convergence rate strongly depends on the Debye length, but also on the initial profile $U^{0}$ (in opposition to the classical Poisson–Nernst–Planck problem without volume filling [3]).
@@ 2.1. Space and time discretizations
For any $K\in\mathcal{T}$, we assume that there exists a subset $\mathcal{E}_{K}$ of distinct elements of $\mathcal{E}$ such that the boundary of a control volume can be described by $\partial K=\bigcup_{\sigma\in\mathcal{E}_{K}}\sigma$ and, consequently, $\mathcal{E}=\bigcup_{K\in\mathcal{T}}\mathcal{E}_{K}$. Additionally, we assume that boundary edges $\mathcal{E}_{\text{ext}}=\mathcal{E}\setminus\mathcal{E}_{\rm int}$ are either subsets of $\Gamma^{D}$ or $\Gamma^{N}$. We denote by $\mathcal{E}^{D},\mathcal{E}^{N}\subset\mathcal{E}$ the sets of boundary edges lying on $\Gamma^{D}$ and $\Gamma^{N}$, respectively. To each control volume $K\in\mathcal{T}$ we assign a cell center ${x}_{K}\in K$ which satisfies the orthogonality condition: If $K,L$ share a face $\sigma=K|L$, then the vector $x_{L}-x_{K}$ is orthogonal to $\sigma=K|L$. The triplet $\left(\mathcal{T},\mathcal{E},\{{x}_{K}\}_{K\in\mathcal{T}}\right)$ is called an admissible mesh.
@@
We finally introduce $\Delta t=\sup_{n\in\mathbb{N}\setminus\{0\}}\tau^{n}$, which we assume to be finite. As we consider the unbounded time interval $\mathbb{R}_{\geq 0}$, we furthermore have to assume that $\inf_{n\geq 1}\tau^{n}>0.$
@@ 2.3. Main results and organization of the paper
$\phi_{\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}\phi\quad\text{and}\quad u_{0,\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}u_{0}\quad\text{in}\;L^{p}_{\rm loc}(\mathbb{R}_{\geq 0};L^{p}(\Omega))\quad\forall p\in[1,+\infty),$
@@
$U_{\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}U\quad\text{in the }L^{\infty}(\mathbb{R}_{\geq 0}\times\Omega)^{I}\text{ weak-$\star$ sense}.$
@@ 3.1. Uniform a priori bounds and existence of a discrete solution
$\mathcal{S}\;:\;\begin{cases}[0,\tau^{n}]\times[-2,2]^{I\times\mathcal{T}}\times[-C-1,C+1]^{\mathcal{T}}\to\mathbb{R}^{\mathcal{T}}\times\mathbb{R}^{\mathcal{T}}\\ (s,U,\phi)\mapsto\mathcal{S}(s,U,\phi)\end{cases}$
@@
Thus, the topological degree corresponding to $\mathcal{S}(s,U,\phi)=0$ is equal to one, all along the homotopy, and hence in particular for $s=\tau^{n}$. That implies the existence of a solution to the scheme (but it does not guarantee uniqueness). ∎
@@ 3.2. Energy dissipation at the discrete level
(41) $\mathcal{H}_{\mathcal{T}}^{n}=\sum_{K\in\mathcal{T}}m_{K}H(U_{K}^{n})+\frac{\lambda^{2}}{2}\sum_{\sigma\in\mathcal{E}}a_{\sigma}(\phi_{K}^{n}-\phi_{K\sigma}^{n})^{2}-\lambda^{2}\sum_{\sigma\in\mathcal{E}^{D}}a_{\sigma}\phi_{\sigma}^{D}(\phi_{\sigma}^{D}-\phi_{K}^{n}),$
@@
where $\widetilde{H}:\left\{(u_{i})_{0\leq i\leq I}\;\middle|\;\sum_{i=0}^{I}u_{i}=1\right\}\to[-\log(I+1),+\infty)$, defined by
$\widetilde{H}(u_{0},\dots,u_{I}):=\sum_{i=0}^{I}u_{i}\log(u_{i}),$
and
@@ 3.3. Further uniform estimates on the discrete solution
$\mathcal{D}_{i,\sigma}^{n}\geq\kappa a_{\sigma}\cosh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right)\left(\sqrt{u_{i,K}^{n}u_{0,L}^{n}}-\sqrt{u_{i,L}^{n}u_{0,K}^{n}}\right)^{2}\\ -2\kappa a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,L}^{n}}+\sqrt{u_{i,L}^{n}u_{0,K}^{n}}\right)^{2}\sinh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right).$
@@
(55) $\sum_{n=1}^{N}\sum_{K\in\mathcal{T}}m_{K}(u_{i,K}^{n}-u_{i,K}^{n-1})\varphi_{K}^{n}\leq C\left(1+t^{N}\right)^{1/2}\left(\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\varphi_{K}^{n}-\varphi_{K\sigma}^{n}\right)^{2}\right)^{1/2}.$